OpenStudy (anonymous):

Find the point on the curve xy=12 that is closest to the point (5,0)

OpenStudy (calculusfunctions):

@nikohellas ready?

OpenStudy (anonymous):

born ready

OpenStudy (anonymous):

born ready

OpenStudy (calculusfunctions):

Alright then. First, do you have any calculus background?

OpenStudy (calculusfunctions):

Okay, so then you're familiar with optimization problems, correct?

OpenStudy (anonymous):

y= 12/x ;d^2 = (x-5)^2 + (12/x)^2

OpenStudy (anonymous):

just cant figure out f'(x).

OpenStudy (calculusfunctions):

Yes, very good! Now for the next step, differentiate the distance d with respect to x, and show me what you get.

OpenStudy (calculusfunctions):

d (distance) is your f(x) here.

OpenStudy (calculusfunctions):

Do you know the 'Power of a Function' rule (chain rule)?

OpenStudy (anonymous):

yes...f' = 2(x-5)(1) + 2(12/x) ?

OpenStudy (calculusfunctions):

Almost. The derivative of 12/x is incorrect. Do you know why?

OpenStudy (calculusfunctions):

It's missing a factor.

OpenStudy (anonymous):

not really.. i get lost when x is the denominator.

OpenStudy (calculusfunctions):

OOPS! Sorry that should be\[f(x)=(x -5)^{2}+(12x ^{-1})^{2}\]

OpenStudy (anonymous):

so f' = 2(x-5)(1) + 2(12X^-1)(-12X^-2)

OpenStudy (calculusfunctions):

Yes! Now first, let's rewrite this derivative with positive exponents. Go ahead!

OpenStudy (anonymous):

2X-10+2(-144/X^3)

OpenStudy (calculusfunctions):

Let's clean it up a bit and leave it as\[f \prime (x)=2(x -5)-\frac{ 288 }{ x ^{2} }\]Correct?

OpenStudy (anonymous):

is that x^2 or x^3 ?

OpenStudy (calculusfunctions):

x squared.

OpenStudy (anonymous):

shouldnt it be x cubed?

OpenStudy (calculusfunctions):

Oh yes, sorry...I'm a bit sleepy still. You're right. It is x cubed\[f \prime (x)=(x -5)^{2}-\frac{ 288 }{ x ^{3} }\]

OpenStudy (anonymous):

so 2(x-5) - 288/x^3

OpenStudy (calculusfunctions):

Again. I meant\[f \prime (x)=2(x -5)^{2}-\frac{ 288 }{ x ^{3} }\]

OpenStudy (anonymous):

where the square power come from?

OpenStudy (calculusfunctions):

Oh my, I'm more tired than I thought. Sorry\[f \prime (x)=2(x -5)-\frac{ 288 }{ x ^{3} }\]

OpenStudy (anonymous):

as am i my friend.. so whats the next step?

OpenStudy (calculusfunctions):

For optimization problems, how do you treat the derivative, to determine the critical numbers?

OpenStudy (anonymous):

Y'=0 ?

OpenStudy (calculusfunctions):

Though technically, critical numbers are the x-values where the first derivative equals zero or where f '(x) is undefined, however for these f '(x) equals zero will suffice. Yes. Go ahead and solve for x.

OpenStudy (anonymous):

x^3(2x-10) = 288

OpenStudy (calculusfunctions):

How about we rewrite it as \[2x ^{3}(x -5)=288\]Divide both sides by 2. Thus\[x ^{3}(x -5)=144\]now expand, move the 144 to the left side, and obtain\[x ^{4}-5x ^{3}-144=0\]OK?

OpenStudy (anonymous):

yes.. ok.. how can i factor that now?

OpenStudy (calculusfunctions):

Well, that's a good question. There is a real root between -4 and -3, and another between 5 and 6. There are no integer or rational roots here. There are irrational roots. Have you learned Newton's method or the intermediate value theorem or the bisection method? If not, than you would use a graphing calculator or an algebra program to solve for the real roots of this equation.

OpenStudy (anonymous):

IVT..

OpenStudy (anonymous):

f(0)<0 and f(2)>0

OpenStudy (calculusfunctions):

Alright, then use the intermediate value theorem (IVT) to determine the irrational roots. To the nearest ten thousandth should be fine.

OpenStudy (anonymous):

im lost again.

OpenStudy (calculusfunctions):

BTW f '(2) < 0. You said f(2) > 0

OpenStudy (calculusfunctions):

You don't remember, how to apply the intermediate value theorem?

OpenStudy (anonymous):

no.

OpenStudy (ybarrap):

I do notice that \(\bf x^4 - 144 = (x^2 - 12)(x^2 + 12)=5x^3\). But it's not clear what next.

OpenStudy (calculusfunctions):

OK. The intermediate value theorem (IVT) states that if a function is continuous on a closed interval a ≤ x ≤ b and there is a value N between f(a) and f(b), then there will be a number x = c in the open interval a < x < b such that f(c) = N. Do you understand as such?

OpenStudy (anonymous):

the answer is (5.76, 2).. how do i get that from the step we're at?

OpenStudy (calculusfunctions):

Did you get that answer by IVT or are you simply stating the given answer?

OpenStudy (calculusfunctions):

Do you understand the theorem, I explained, above? Yes or no?

OpenStudy (calculusfunctions):

Alright, using this equation, let's work through the theorem. Ready?

OpenStudy (calculusfunctions):

I see you closed the question, without responding to my questions or even a mention of closing. How rude of you to waste someone's time like that. @nikohellas

OpenStudy (ybarrap):

I'd be interested to know how you are thinking of proceeding?

OpenStudy (calculusfunctions):

@ybarrap of course.

OpenStudy (ybarrap):

Sounds like you are thinking of using a numerical technique like newton's method

OpenStudy (calculusfunctions):

First, do you understand the theorem, I explained, above?

OpenStudy (ybarrap):

i was trying to depress the factors the polynomial

OpenStudy (ybarrap):

yes, I'm familiar with it

OpenStudy (calculusfunctions):

No, this is a prime polynomial (cannot be factored over integers).

OpenStudy (ybarrap):

is there a test for that?

OpenStudy (ybarrap):

rational root theorem?

OpenStudy (calculusfunctions):

OK. Let's begin. Let our derivative function be g(x). First, determine two values x = a and x = b such that g(a) < 0 and g(b) > 0 or g(a) > 0 and g(b) < 0. No the rational root theorem will not work because there are no rational roots either.

OpenStudy (ybarrap):

ok

OpenStudy (ybarrap):

g(1) < 0 and g(10) > 0

OpenStudy (calculusfunctions):

Try g(5) and g(6) for example. To the nearest thousandth, tell me the value of each. Go ahead please.

OpenStudy (calculusfunctions):

Sorry my server is giving me problems again.

OpenStudy (ybarrap):

-144,72

OpenStudy (ybarrap):

This is our function, right? \(\bf g(x)=x^4-5x^3-144\)

OpenStudy (ybarrap):

g(5)=-144,g(6)=72 g(-3)=72,g(-2)=-88

OpenStudy (calculusfunctions):

g(6) =?

OpenStudy (ybarrap):

72

OpenStudy (ybarrap):

g(6)=72

OpenStudy (ybarrap):

So there is an extrema between x=5 an x=6

OpenStudy (calculusfunctions):

Yes

OpenStudy (calculusfunctions):

Now since g(6) is closer to zero, than g(5), give me g(5.6) to the nearest thousandth.

OpenStudy (ybarrap):

4.95

OpenStudy (ybarrap):

*-38.63

OpenStudy (ybarrap):

g(5.753)=-.63

OpenStudy (ybarrap):

g(5.7553)=-0.01

OpenStudy (calculusfunctions):

Then try g(5.7) and g(5.8). You will notice that g(5.7) is negative and g(5.8) is positive. However, g(5.8) is closer to zero than g(5.7). Hence try a value of x closer to 5.8, between 5.7 and 5.8. Continuing this process, you should get very close to zero.

OpenStudy (ybarrap):

So and poor-man's newton's method. I sometimes forget that brute-force has its place. Thanks!

OpenStudy (ybarrap):

x = 5.7553

OpenStudy (ybarrap):

Wolfram's positive root solution: $$\tt x=\dfrac{1}{4} \sqrt{8 \sqrt[3]{3 \left(\sqrt{17913}-75\right)}-\dfrac{128\ 3^{2/3}}{\sqrt[3]{\sqrt{17913}-75}}+25}+\\\dfrac{1}{2} \sqrt{-2 \sqrt[3]{3 \left(\sqrt{17913}-75\right)}+\dfrac{32\ 3^{2/3}}{\sqrt[3]{\sqrt{17913}-75}}+\\\dfrac{125}{2 \sqrt{8 \sqrt[3]{3 \left(\sqrt{17913}-75\right)}-\dfrac{128\ 3^{2/3}}{\sqrt[3]{\sqrt{17913}-75}}+25}}+\dfrac{25}{2}}+\dfrac{5}{4}$$ (and I was trying to depress the quartic) \(\Large\ddot\smile\)

OpenStudy (calculusfunctions):

Yes, like the quadratic formula, there are formulas to solve cubic and quartic equations, but they're very tedious so we don't put them in practice.

OpenStudy (ybarrap):

Thank goodness!